# Data Mesh: a Systematic Gray Literature Review

### v1

@@ 5.1. A Layered Architecture of Capabilities and Roles in Data Mesh
We identified six different roles for data mesh actors (see Table 7):
@@
are responsible for building and maintaining the self-serve platform. They also create the IaC blueprints that product developers can use to provision and configure infrastructure resources and platform services for hosting and managing data products.
@@
are responsible for offering and governing the data products in their domains. They must ensure that the data products are interoperable and meet the requirements of the consumers.
@@
build and publish data products on the data mesh using the tools/services the self-serve platform provides.
@@
use data from the data products. Product aggregators are also product consumers since they consume multiple other data products and aggregate them into composite data products.
@@
comprise representatives from different domains across the organization and governs the mesh globally.
Table 7. Roles in data mesh: codes and sources.
Role Frequency Unique Sources
Data Platform Team 7 S2,S22,S42,S73,S85,S86
Product Owner 17 S1,S2,S3,S16,S18,S25,S33,S35,S43,S49,S67,S73,S82,S86-S88
Product Developer 10 S1,S2,S16,S18,S21,S35,S39
Product Consumer 9 S10,S13,S19,S54,S91
Federated Governance Team 9 S1,S16,S21,S25,S28,S110
@@
The data product layer sits on top of the self-serve layer. As discussed in Section 3.1, data products consist of data, metadata, code, interfaces, and the infrastructure to deploy the product. Once the product developers define and implement all components, the data products can be published in a product catalog. Product consumers browse the catalog to discover candidate data products for their use cases and check the corresponding metadata to verify that they suit their needs. Product meta-data can also enable consumers to automatically connect (binding) to the data products to obtain their data.
Composite data products use data from other (basic or composite) data product(s), potentially apply complex transformations, and publish the results as new data products (see Section 3.1). Composite data products contain the same components as basic data products. However, product aggregators perform additional actions to build composite data products, similar to the additional responsibilities in xSOA (Papazoglou and Van Den Heuvel, 2007). To distinguish between these responsibilities, the layered architecture in Figure 7 differentiates between product developers, who build basic data products, and product aggregators, who create composite data products.
Composite data products need to provide new value to the organization. A simple integration of two data sources can easily be replicated and may not provide sufficient additional value. Therefore, a product aggregator may need to perform complex transformations on the data from upstream data products to create value-added data. Furthermore, composing products may involve coordinating a series of steps, e.g., ingesting data from multiple data products, merging the data from a subset set of products, cleaning and standardizing data, and storing intermediate and final results.
Composite data products must conform with the upstream data products throughout their lifetime. For example, upstream product developers might change their data products (e.g., schema and product interfaces), which can affect downstream consumers. Product aggregators must deal with such issues by adopting suitable strategies such as product versioning and using an anti-corruption layer (an adapter layer).
@@
Data products must be managed and governed at runtime to ensure they provide value to the organization. As discussed in Section 3, there are two levels of governance: local and global. The local governance (per each product) is responsible for the quality and conformance of individual data products. For example, the local governance team must ensure that a data product operates within the specified SLOs throughout its lifetime and remains compliant with local and global policies and external regulations. The owners of the composite products also need to monitor the performance of upstream data products in terms of service level objectives (SLOs). The performance of the composite data products can be affected by that of upstream data products. Thus, a product aggregator needs to monitor the performance of upstream products to determine any SLO violations and ensure the composite products operate within its SLOs. The main goal of global governance at runtime is to ensure that the mesh architecture can scale throughout the organization effectively. The federated governance team monitors the mesh to assess the data mesh’s overall health and enforce the global policies, interoperability standards, and guidelines.
@@ 5.2. A Layered Architecture for Data Mesh Development
The data mesh literature does not explicitly provide an end-to-end methodology for implementing data mesh in or across organizations. However, the literature advocates applying the principles of domain-driven design (DDD), product thinking, and platform-based development. In particular, data products should be identified by applying DDD to the data produced by the business domains of the organization (see Table 8). Inspired by a layered model for SOA development (Papazoglou, 2008), we introduce a layered model for data mesh development that integrates the previously mentioned principles. Figure 8 shows the proposed layered approach to data mesh development. The model comprises seven distinct layers: business domains, data domains, data-bounded contexts, data products, data product components, self-serve platform services, and operating infrastructure.
Table 8. Applying domain-driven design to develop data products: sources and code frequency.
Concern Frequency Unique Sources
Applying Domain-Driven Design 28
S1,S2,S11,S36,S37,S39,S45,S46,S53,S55,S58,S71,S75,
S78,S79,S87,S90,S91,S96,S98,S103,S110,S111,S112
Business Domain and Data Domain Alignment 14
S36,S37,S39,S46,S53,S55,S56,S71,S75,
S78, S87, S111, S112
Bounded Contexts for Data Products 14 S1,S35,S45,S79,S87,S90,S98,S112
@@
Following the DDD method, the first layer decomposes an organization’s business into disjoint domains. A business domain consists of a set of business processes that shares a common set of business capabilities and collaborate to accomplish the business goals and objectives of the domain (Papazoglou, 2008). These processes produce valuable data for the domain and the rest of the organization and can thus be offered as data products. For example, a manufacturing organization may include domains such as distribution, manufacturing, finance, and human resources (Papazoglou, 2008). The distribution domain may comprise key business processes such as consumer purchasing, order management, and inventory. The purchasing process may produce data containing information about products or services, time of purchase, and the amount spent. This data may be exploited to build data products to answer queries about purchase history, customer buying patterns, and other relevant details, such as stock availability and product appearance.
Data Domains.
@@
The second layer partitions the analytical data of a business domain into one or more data domains. Data domains define boundaries around the analytical data. From the viewpoints of data mesh and data governance, data domains categorize and organize data to assign accountability and responsibility for the data. As discussed above, different business processes in a business domain may produce multiple data. Due to the diversity and complexity of processes, the responsibility for each process may be assigned to a separate team (Danilova, 2019). Similarly, the analytical data of processes may be owned by different teams, originating multiple data domains from the same business domain. The data mesh gray literature also recommends aligning business and data domains but creating new domains for shared data as appropriate.
Data-bounded Contexts.
@@
The third layer identifies the bounded contexts within each data domain. DDD designs and implements software systems based on domain models, which are the object models of the domains that incorporate both behavior and data in the domain (Fowler, 2012; Evans and Evans, 2004). Creating and maintaining a unified model for larger and more complex domains, where multiple teams work on separate subsystems, may not be feasible or cost-effective (Fowler, 2012; Evans and Evans, 2004). To tackle this issue, DDD partitions the solution space of a domain into bounded contexts, where each can have a separate domain model. A data domain can also be complex and include multiple logical groupings of data. Each group can have a separate data owner and model (i.e., bounded contexts for decomposing data domains). For example, there may be different categories of customers: online only, physical only, third party, prospect, individual, group, corporation, government, and charity. Managing data for each or subset of these categories might differ, and multiple teams or individuals might be responsible for it. Consequently, multiple bounded contexts emerge within a customer data domain.
@@
Various dependencies (design time and runtime) may exist between data bounded contexts. The gray literature recommends using the strategic design patterns of DDD (e.g., customer-supplier and anti-corruption layer) to integrate bounded contexts (i.e., context mapping). For example, a bounded context of a composite product can use the anti-corruption layer to minimize the impacts of the changes to the data models of the data products from the upstream bounded contexts.
Data Products.
@@
The fourth layer identifies data products within each bounded context. For building applications for a bounded context, DDD recommends applying so-called tactical design patterns: entity, value object, aggregate, service, repository, factory, event, and module. These patterns can be used to enrich a domain model of a bounded context, and the enriched model can be used to derive the application architecture components. For example, in a microservice architecture, services, APIs, and events can be deduced from the domain model using the tactical design patterns (Singjai et al., 2021). The data mesh gray literature also recommends applying these patterns, for example, using domain main events for ingesting and serving data and implementing a log of analytical data changes, and using entities and aggregates for identifying products and creating database schemas and product APIs.
Data Products Components.
@@
The fifth layer consists of the components that implement functional and non-functional requirements of data products. There are different types of data products, for example, datasets (offered as database views or APIs), data pipelines, ML training pipelines, and ML prediction services. Different data products may use different architectures. For example, data pipeline products may use popular data architecture styles such as Lambda and Kappa (Davoudian and Liu, 2020), and ML prediction services may use architecture styles as model-as-a-service or model-as-a-dependency (Warnett and Zdun, 2022). The resulting data products can have components like streaming data ingestion, ETL workflows, storage of raw, cleaned, and enriched data, model and feature storage, APIs, data quality assessors, and report generators. Data products can also share the implementation of some components, for example, a generic data validating service or model validation service.
Self-serve Platform Services.
@@
The sixth layer consists of the self-serve platform services that support the end-to-end lifecycle of data products and their components, from design, implementation, testing, deployment, execution, and monitoring, to evolution. For example, data product developers should be able to use an ETL platform service in a self-service manner to create data ELT pipelines for their data products, deploy and execute the created pipelines on-demand or according to a schedule, monitor the performance, resource usage, and errors of the executed pipelines, and make changes to the pipeline as necessary. Section 3.4 lists common platform services we found from the gray literature. The platform team is responsible for self-serve platform services. They may develop platform services from scratch, reuse and customize third-party software, use the platform services offered by cloud providers, and modernize existing platform services by incorporating self-service capabilities.
Operating Infrastructure.
@@ 5.3. A Logical Runtime Structure of Data Mesh
This section presents a logical runtime view of a data mesh. The data mesh architectures presented in the gray literature source primarily focus on the runtime model of an inter-connected web of managed (governed) data products across data domains (see Table 9). We extracted the essential components from those architecture designs and created a consolidated model by consulting web service management architectures (Papazoglou et al., 2007; Papazoglou, 2008; Li et al., 2019; Zhou et al., 2002; Kumara et al., 2017). Figure 9 depicts the reference architecture at runtime for data mesh.
@@
Table 9. Logical runtime architecture: codes and sources.
Sub Category Atomic Codes
ID Name Frequency ID Name Frequency Unique Sources
6.1
Overall
Structure
21 6.1.1
Logical Structure with Three Layers
(Self-serve Platform, Products, and Governance)
21
S1,S8-S10,S14,S29,S35,S42,S47,S53,
S58,S70,S71,S78,S80,S86,S99,S110
6.2
Product
Interfaces
25 6.2.1 Input 9 S51,S66,S72,S86,S97,S103,S108,S110
6.2.2 Output 9 S51,S66,S72,S86,S97,S103,S108,S110
6.2.3 Management 7 S86,S97,S103,S110
6.3
Communication
Channels
40 6.3.1 API 12 S1,S64,S71,S75,S78,S82,S83,S90,S108,S111
6.3.2 Message/Event Bus 18 S24,S29,S64,S78,S82,S83,S84,S86,S90,S108,S110,S111
6.3.3 Shared Storage 10 S13,S24,S86,S87,S89,S97,S98,S100,S112,S108
6.4
Management
Agents
13 6.4.1 Sidecar Proxy 9 S29,S36,S78,S83,S113
6.4.2 Policy Agent 4 S2,S29,S78,S97
@@
A product container is a deployment and runtime environment for a data product and its managing agents. We borrowed the concept of a container and a management agent from Web services containers and management agents (Papazoglou, 2008). As in the web service management architectures (Papazoglou, 2008; Papazoglou et al., 2007; Li et al., 2019), we separate the communication between data products, source systems, and end-users (data channel and input/output interfaces) and the communication between the management agents and the management/governance plane (management channel and management interface). The input interface enables data products to receive data from source systems and other data products. The output interface allows data products to provide their data to consumers or composite data products. Finally, the management interfaces enable monitoring and controlling the behaviors of data products (e.g., monitor product performance, raising alerts, controlling data routing, and enforcing security policies and data contracts). The modes of communication can be pull-based (e.g., retrieving data via a REST API or a database connector or polling data from a message broker) or push-based (e.g., pushing the data to the downstream consumers as the data become available).
@@
The management agents of a product implement management tasks at the product level, such as marking and enforcing policy decisions, collecting performance metrics, and managing the product life cycle (i.e., creating, updating, and destroying products). The gray literature recommends using the sidecar design pattern and the policy-as-code approach to implement the monitoring and policy enforcement capabilities of the management agents. The sidecar pattern (Burns and Oppenheimer, 2016) decouples policy decision-making from the data product and can enable reuse and consistent implementation of monitoring and management logic across data products. The policy-as-code approach uses programming code to write policies, allowing applying software development best practices such as version control, automated testing, and automated deployment. A policy manifests as a set of rules that dictate the behaviors of the management agents, for example, how to enforce authentication and authorization policies for data product access, when to raise alerts, and where to route messages.
@@
The components/applications in the management and governance plane can use the management interfaces of products to receive information (e.g., performance data, access logs, and alerts) from data products and to send configuration and policy updates to the data products. We identified several key components for this plane from the gray literature: data product registry and discovery service, policy authoring service, policy repository, global metadata repository, and data mesh dashboard. The product registry and discovery service is connected to the management channel. Product developers publish their products to it over the management channel. Consumers can discover data products by using the published metadata of data products. The federated governance team can create global policies (e.g., interoperability, security, and compliance) using the policy authoring services, store them in the policy repository for reuse and versioning, and enact them using the management interfaces provided by the individual products. The team can use the data mesh dashboard to monitor the operational health of the data mesh, including the cross-domain access and usage of data products and level of policy compliance.
@@
The self-serve data platform plane provide infrastructure resources and platform services to develop, test, build, deploy, execute, monitor, and manage data products and governance components. For example, a data product can use the ML workflow platform service to create and run the training pipelines. The product registry and discovery service can use data/product catalog or metadata repository platform services. Section 3.4 discusses the key platform services of a self-serve data platform.
@@
In this section, we put forward a set of research challenges for data mesh that need to be addressed. We identified them by mapping the challenges mentioned by the practitioners in the gray literature to the relevant research issues from the service-oriented computing (Papazoglou et al., 2007; Wei and Blake, 2010; Papazoglou and Van Den Heuvel, 2007) and data management (Dustdar et al., 2012; Stonebraker et al., 2018; Al-Ruithe et al., 2019; Janssen et al., 2020a). Table 10 shows the challenges proposed by the practitioners.
Table 10. Data mesh challenges: sources and code frequencies.
Challenge Code Frequency Unique Sources
Standardizing Data Mesh 6 S8,S35,S38,S61,S110,S111
Methodologies and Tools for Data Mesh Development and Operation 18 S1,S4,S28,S29,S35,S42,S98,S111
Data Product Life Cycle 14 S1,S35,S45,S79,S87,S90,S98,S112
Self-serve Platform Services 7 S2,S16,S33,S43,S53,S86,S108
Data Mesh Governance 7 S1,S28,S35,S86,S111
Organizational Change Management 8 S2,S16,S28,S33,S35,S61,S87

### v2

@@ 5.1. A Layered Architecture of Capabilities and Roles in Data Mesh
As discussed in Section 3.1, Zhamak and the gray literature categorize data products into source-aligned, aggregates, and consumer-aligned. By adopting service types in SOA, we classify them into atomic and composite data products. Like composite web services, composite data products use other data products (atomic or composite) for their source data and integrate them to create new value-added products. Hence, aggregates and consumer-aligned products are composite products. The source-aligned data products are atomic because they only ingest data from the operational source systems.
@@
[…]
The layered architecture of capabilities and roles in data mesh.
@@
We identified six different roles for data mesh actors:
@@
are responsible for building and maintaining the self-serve data platform (Dehghani, 2019b; Mitrus, 2021; Oostra, 2021b; Raab and Chiapusso, 2021; Moses, 2022; Kumar, 2022). They also create the IaC blackprints that product developers can use to provision and configure infrastructure resources and platform services for hosting and managing data products. They need to possess software and data engineering skills (Mitrus, 2021).
@@
are responsible for offering and governing the data products in their domains (Dehghani, 2022b, 2019b; Jarvin Mutatiina and Blaauw, 2020; Gillin, 2021; ELEKS, 2021; Thandavarayan, 2021; Keboola, 2021; Rajagopalan, 2021; Snowflake, 2021; Sequeda, 2021; Pungas, 2021; Raab and Chiapusso, 2021; Broda, 2022a; Kumar, 2022; Broda, 2022c; Thoughtworks, 2021). They must ensure that the data products are interoperable and meet the requirements of the consumers. They are generally responsible for domain-level governance activities (see Section 3.3).
@@
build and publish data products on the data mesh using the tools/services the self-serve platform provides (Dehghani, 2022b, 2019b; Gillin, 2021; ELEKS, 2021; Riccomini, 2021; Rajagopalan, 2021; Das, 2020).
@@
use data from the data products (JPMorgan, 2021; Nivas Shankar and Hasson, 2021; Lam, 2021; Strengholt, 2021; Abedrabbo, 2021). Product aggregators are also product consumers since they consume multiple other data products and aggregate them into composite data products.
@@
comprise representatives from different domains across the organization and govern the mesh globally (Dehghani, 2022b; Gillin, 2021; Riccomini, 2021; Thandavarayan, 2021; Perlov, 2021; Jochen Christ and Harrer, 2022). Section 3.3 discusses the governance activities of the federated governance team.
@@
The data product layer sits on top of the self-serve layer. As discussed in Section 3.1, data products consist of data, metadata, code, interfaces, and the infrastructure to deploy the product. Once the product developers define and implement all components, the data products can be published in a product catalog. Product consumers browse the catalog to discover candidate data products for their use cases and check the corresponding metadata to verify that they suit their needs. Product metadata can also enable consumers to automatically connect (binding) to the data products to obtain their data. For example, if a data product uses a storage bucket as a serving endpoint, it can publish the metadata, such as the URL of the bucket and the names and schemas of the result objects, to the catalog (Orwat, 2021).
Composite data products use data from other (atomic or composite) data product(s), potentially apply complex transformations, and publish the results as new data products (see Section 3.1). Composite data products contain the same components as atomic data products. However, product aggregators perform additional actions to build composite data products, similar to the additional responsibilities in xSOA (Papazoglou and Van Den Heuvel, 2007). To distinguish between these responsibilities, the layered architecture in Figure 6 differentiates between product developers, who build atomic data products, and product aggregators, who create composite data products.
Composite data products need to provide new value to the organization. A simple integration of two data sources can easily be replicated and may not provide sufficient additional value. Therefore, a product aggregator may need to perform complex transformations on the data from upstream data products to create value-added data. Furthermore, composing products may involve coordinating a series of steps,e.g.,, ingesting data from multiple data products, merging the data from a subset set of products, cleaning and standardizing data, and validating and storing intermediate and final results.
Composite data products must conform with the upstream data products throughout their lifetime. For example, upstream product developers might change their data products (e.g., schema and product interfaces), which can affect downstream consumers (Cooke, 2021; Stirling, 2021; Baker, 2021a; Jochen Christ and Harrer, 2022, 2022). Product aggregators must deal with such issues by adopting suitable strategies such as product versioning and using an anti-corruption layer (an adapter layer) (Baker, 2021a; Jochen Christ and Harrer, 2022).
@@
Data products must be managed and governed at runtime to ensure they provide value to the organization. As discussed in Section 3.3, there are two levels of governance: local and global. The local governance (per each product) is responsible for the quality and conformance of individual data products. For example, the local governance team must ensure that a data product operates within the specified SLOs throughout its lifetime and remains compliant with local and global policies and external regulations. The owners of the composite products also need to monitor the performance of upstream data products in terms of service level objectives (SLOs). The performance of the composite data products can be affected by that of upstream data products. Thus, a product aggregator needs to monitor the performance of upstream products to determine any SLO violations and ensure the composite products operate within its SLOs. The main goal of global governance at runtime is to ensure that the mesh architecture can scale throughout the organization effectively. The federated governance team monitors the mesh to assess the data mesh’s overall health and enforce the global policies, interoperability standards, and guidelines.
@@ 5.2. A Layered Architecture for Data Mesh Development
The data mesh literature does not explicitly provide an end-to-end methodology for implementing data mesh in or across organizations. However, the literature advocates applying the principles of domain-driven design (DDD), product thinking, and platform-based development. In particular, data products should be identified by applying DDD to the data produced by the business domains of the organization (Dehghani, 2022b, 2019b; Pollock, 2021; Dehghani, 2021b; Limited, 2021; Das, 2020; Orwat, 2021; Engineering, 2021; Cooke, 2021; Stirling, 2021; Moses, 2021; Baker, 2021b; McKenzie, 2021; Oracle, 2021b; Reid, 2021; Strenholt and Courtright, 2022; Strengholt, 2020; Abedrabbo, 2021; Platter, 2021b; Ramos, 2021; Maharana, 2021; Jochen Christ and Harrer, 2022; Strengholt, 2022c, b). Inspired by a layered model for SOA development (Papazoglou, 2008), we introduce a layered model for data mesh development that integrates the previously mentioned principles. Figure 7 shows the proposed layered approach to data mesh development. The model comprises seven distinct layers: business domains, data domains, data-bounded contexts, data products, data product components, self-serve platform services, and operating infrastructure.
[…]
A layered model for the data mesh development.
@@
Following the DDD method, the first layer decomposes an organization’s business into disjoint domains. A business domain consists of a set of business processes that share a common set of business capabilities and collaborate to accomplish the business goals and objectives of the domain (Papazoglou, 2008). These processes produce valuable data for the domain and the rest of the organization and can thus be offered as data products. For example, a manufacturing organization may include domains such as distribution, manufacturing, finance, and human resources (Papazoglou, 2008). The distribution domain may comprise key business processes such as consumer purchasing, order management, and inventory. The purchasing process may produce data containing information about products or services, time of purchase, and the amount spent. This data may be exploited to build data products to answer queries about purchase history, customer buying patterns, and other relevant details, such as stock availability and product appearance.
@@
The second layer partitions the analytical data of a business domain into one or more data domains. Data domains define boundaries around the analytical data. From the viewpoints of data mesh and data governance, data domains categorize and organize data to assign accountability and responsibility for the data. As discussed above, different business processes in a business domain may produce multiple data. Due to the diversity and complexity of processes, the responsibility for each process may be assigned to a separate team (Danilova, 2019). Similarly, the analytical data of processes may be owned by different teams, originating multiple data domains from the same business domain. The gray literature also recommends aligning business and data domains and creating new domains for shared data as appropriate (Dehghani, 2021b; Limited, 2021; Das, 2020; Engineering, 2021; Cooke, 2021; Stirling, 2021; Rasool, 2021; Baker, 2021b; McKenzie, 2021; Oracle, 2021b; Strenholt and Courtright, 2022; Strengholt, 2022c, b).
@@
The third layer identifies the bounded contexts within each data domain. The gray literature recommends using bounded contexts for scoping data products (Dehghani, 2022b; Rajagopalan, 2021; Orwat, 2021; Reid, 2021; Strenholt and Courtright, 2022; Strengholt, 2020; Ramos, 2021; Strengholt, 2022b). DDD designs and implements software systems based on domain models, which are the object models of the domains that incorporate both behavior and data in the domain (Fowler, 2012; Evans and Evans, 2004). Creating and maintaining a unified model for larger and more complex domains, where multiple teams work on separate subsystems, may not be feasible or cost-effective (Fowler, 2012; Evans and Evans, 2004). To tackle this issue, DDD partitions the solution space of a domain into bounded contexts, where each can have a separate domain model. A data domain can also be complex and include multiple logical groupings of data. Each group can have a separate data owner and model (i.e., bounded contexts for decomposing data domains). For example, there may be different categories of customers: online only, physical only, third party, prospect, individual, group, corporation, government, and charity. Managing data for each or subset of these categories might differ, and multiple teams or individuals might be responsible for it. Consequently, multiple bounded contexts emerge within a customer data domain.
@@
Various dependencies (design time and runtime) may exist between data bounded contexts. The gray literature recommends using the strategic design patterns of DDD (e.g., customer-supplier and anti-corruption layer) to integrate bounded contexts (i.e., context mapping) (Strengholt, 2022c). For example, a bounded context of a composite product can use the anti-corruption layer to minimize the impacts of the changes to the data models of the data products from the upstream bounded contexts (Jochen Christ and Harrer, 2022). For example, a data product for sales funnel analysis uses the customer search data product to get customers’ browsing history. The funnel analysis product can implement an anti-corruption layer to prevent the adverse effect of changes in browsing history data format.
@@
The fourth layer identifies data products within each bounded context. For building applications for a bounded context, DDD recommends applying so-called tactical design patterns: entity, value object, aggregate, service, repository, factory, event, and module. These patterns can be used to enrich a domain model of a bounded context, and the enriched model can be used to derive the application architecture components. For example, in a microservice architecture, services, APIs, and events can be deduced from the domain model using the tactical design patterns (Singjai et al., 2021). The data mesh gray literature (Baker, 2021b; Platter, 2021b; Maharana, 2021; Jochen Christ and Harrer, 2022) also recommends applying these patterns, for example, using domain main events for ingesting and serving data and implementing a log of data changes, and using entities and aggregates for identifying products and creating database schemas and product APIs.
@@
The fifth layer consists of the components that implement functional and non-functional requirements of data products. Different types of data products exist, such as datasets (offered as database views or APIs), data pipelines, ML training pipelines, and ML prediction services. Different data products may use different architectures. For example, data pipeline products may use popular data architecture styles such as Lambda and Kappa (Davoudian and Liu, 2020), and ML prediction services may use architecture styles as model-as-a-service or model-as-a-dependency (Warnett and Zdun, 2022). The resulting data products can have components like streaming data ingestion, ETL workflows, storage of raw, cleaned, and enriched data, model and feature storage, APIs, data quality assessors, and report generators. Data products can also share the implementation of some components, such as a generic data validating service or model validation service.
@@
The sixth layer consists of the self-serve platform services that support the end-to-end lifecycle of data products and their components, from design, implementation, testing, deployment, execution, and monitoring to evolution. For example, data product developers should be able to use an ETL platform service in a self-service manner to create ETL pipelines for their data products, deploy and execute the created pipelines on-demand or according to a schedule, monitor the performance, resource usage, and errors of the executed pipelines, and make changes to the pipeline as necessary. Section 3.4 lists common platform services we found from the gray literature. The platform team is responsible for self-serve platform services. They may develop platform services from scratch, reuse and customize third-party software, use the platform services offered by cloud providers, and modernize existing platform services by incorporating self-service capabilities.
@@ 5.3. A Logical Runtime Structure of Data Mesh
This section presents a logical runtime view of a data mesh. The data mesh architectures presented in the gray literature source primarily focus on the runtime model of an inter-connected web of managed (governed) data products across data domains (Dehghani, 2022b; AgileLab, 2021; Furia, 2021; JPMorgan, 2021; Vellante, 2021; Oracle, 2021a; Rajagopalan, 2021; Oostra, 2021b; Dalvik, 2020; Cooke, 2021; Moses, 2021; Ericsson, 2021; Baker, 2021b; Oracle, 2021b; Wells, 2021; Kumar, 2022; Ramos, 2022; Velante, 2021; Jochen Christ and Harrer, 2022). We extracted the essential components from those architecture designs and created a consolidated model by consulting web service management architectures (Papazoglou et al., 2007; Papazoglou, 2008; Li et al., 2019; Kumara et al., 2017). Figure 8 depicts the reference architecture at runtime.
[…]
Logical runtime structure of a data mesh.
@@
A product container is a deployment and runtime environment for a data product and its managing agents. We borrowed the concept of a container and a management agent from Web services containers and management agents (Papazoglou, 2008). As in the web service management architectures (Papazoglou, 2008; Papazoglou et al., 2007; Li et al., 2019), we separate the communication between data products, source systems, and end-users (data channel and input/output interfaces) and the communication between the management agents and the management/governance plane (management channel and management interface). The input interface (Orr, 2020; Adevinta, 2021; Baker, 2021a; Kumar, 2022; Dehghani, 2021a; Maharana, 2021; Wolf, 2022; Jochen Christ and Harrer, 2022) enables data products to receive data from source systems and other data products. The output interface (Orr, 2020; Adevinta, 2021; Baker, 2021a; Kumar, 2022; Dehghani, 2021a; Maharana, 2021; Wolf, 2022; Pollock, 2021) allows data products to provide their data to consumers or composite data products. Finally, the management interfaces (Kumar, 2022; Dehghani, 2021a; Maharana, 2021; Jochen Christ and Harrer, 2022) enable monitoring and controlling the behaviors of data products (e.g., monitor product performance, raising alerts, controlling data routing, and enforcing security policies and data contracts). The modes of communication can be pull-based (e.g., retrieving data via a REST API (Dehghani, 2022b; Madan, 2021; Baker, 2021b; McKenzie, 2021; Oracle, 2021b; Broda, 2022a, b; Strengholt, 2020; Wolf, 2022; Strengholt, 2022c) or reading from a shared database/bucket (Nivas Shankar and Hasson, 2021; Hicks, 2021; Kumar, 2022; Strenholt and Courtright, 2022; Modi, 2021; Dehghani, 2021a; Ramos, 2021; Schultze and Wider, 2020a; Strengholt, 2022b; Wolf, 2022) or polling data from a message broker (Hicks, 2021; Oracle, 2021a; Madan, 2021; Oracle, 2021b; Broda, 2022a, b, d; Kumar, 2022; Strengholt, 2020; Broda, 2022e; Wolf, 2022; Jochen Christ and Harrer, 2022; Strengholt, 2022c)) or push-based (e.g., pushing the data to the downstream consumers as the data become available (Nivas Shankar and Hasson, 2021; Baker, 2021a; Broda, 2022e; Jochen Christ and Harrer, 2022; Dehghani, 2022a)).
@@
The management agents (Dehghani, 2019b; Oracle, 2021a, b; Dehghani, 2021a) of a product implement management tasks at the product level, such as marking and enforcing policy decisions, collecting performance metrics, and managing the product life cycle (i.e., creating, updating, and destroying products). The gray literature recommends using the sidecar design pattern and the policy-as-code approach to implement the monitoring and policy enforcement capabilities of the management agents (Oracle, 2021a; Dehghani, 2021b; Oracle, 2021b; Broda, 2022b; Strengholt, 2022a). The sidecar pattern (Burns and Oppenheimer, 2016) decouples policy decision-making from the data product and can enable reuse and consistent implementation of monitoring and management logic across data products. The policy-as-code approach uses programming code to write policies, allowing applying software development best practices such as version control, automated testing, and automated deployment. A policy manifests as a set of rules that dictate the behaviors of the management agents, for example, how to enforce authentication and authorization policies for data product access, when to raise alerts, and where to route messages.
@@
The components/applications in the management and governance plane can use the management interfaces of products to receive information (e.g., performance data, access logs, and alerts) from data products and to send configuration and policy updates to the data products. We identified several key components for this plane from the gray literature: data product registry and discovery service (Moses, 2020; Broda, 2022c; Thoughtworks, 2021; Ramos, 2021), policy authoring service and policy repository (Cooke, 2021; Jochen Christ and Harrer, 2022), global metadata repository (Dalvik, 2020; Cooke, 2021; Jochen Christ and Harrer, 2022; Strengholt, 2022f), and data mesh dashboard (Cooke, 2021; Jochen Christ and Harrer, 2022). The product registry and discovery service is connected to the management channel. Product developers publish the metadata of their products over the management channel. Consumers can discover data products by using the published metadata. The global metadata repository can store metadata shared across domains, such as master and reference data. The federated governance team can create global policies (e.g., interoperability, security, and compliance) using the policy authoring services, store them in the policy repository for reuse and versioning, and enact them using the management interfaces provided by the individual products. The team can use the data mesh dashboard to monitor the operational health of the data mesh, including the cross-domain access and usage of data products and level of policy compliance.
@@
The self-serve platform plane provides infrastructure resources and platform services to develop, test, build, deploy, execute, monitor, and manage data products and governance components. For example, a data product can use the ML workflow platform service to create and run the training pipelines. The product registry and discovery service can use data/product catalog or metadata repository platform services. Section 3.4 discusses the key platform services of a self-serve platform.
@@
In this section, we put forward a set of research challenges for data mesh that need to be addressed. We identified them by mapping the challenges mentioned by the practitioners in the gray literature to the relevant research issues from the service-oriented computing (Papazoglou et al., 2007; Wei and Blake, 2010; Papazoglou and Van Den Heuvel, 2007) and data management (Dustdar et al., 2012; Stonebraker et al., 2018; Al-Ruithe et al., 2019; Janssen et al., 2020).